OpenStudy (candycane):

Solve the quadratic equation by completing the square: x^2+6x-5=0. I have taken the practice test this is merely for extra credit, please help if at all possible.

OpenStudy (tkhunny):

Well, then give it a go. It will look like this: (x^2 + 6x + _________) - 5 - ___________ You just have to find the magic, missing number. Think long and hard on this: \((a+b)^{2} = a^{2} + 2ab + b^{2}\)

OpenStudy (candycane):

being as I have the answers in front of me I know the missing numbers are -3 and a positive or negativ 14 inside a square root box. I'm not sure how to get to these answers

OpenStudy (tkhunny):

Did you look long and hard at that thing I gave you? The answer is in there.

OpenStudy (candycane):

Yes I looked at it, its rather confusing and I do not understand

OpenStudy (tkhunny):

Okay, let's try this. This is your hint: \((a+b)^{2} = a^{2} + 2ab + b^{2}\) Keep that in mind. A Martian confronts you and gives you a problem: \(a^{2} + 2ab\;+\;...\) The alien invader's demand is that you find the third term that completes the square or Earth will be destroyed. What say you? Shall we live or be vaporized?

OpenStudy (marigirl):

i can try if u like?

OpenStudy (candycane):

I give you credit for creativity but i still dont understand...

OpenStudy (candycane):

Would you please?

OpenStudy (marigirl):

wait, i dunoo if i can... are u meaning from getting from x^2+6x-5 to (x+3)^2-14 .. or beyond that

OpenStudy (tkhunny):

Thus, we lose our lovely planet. Seriously? Let's think about this... 1) I square a simple binomial. \((a+b)^{2} = a^{2} + 2ab + b^{2}\) You say it is rather confusing. I am not allowed to believe this. Are you saying you NEVER have multiplied two binomials? \((a+b)(a+b) = a^{2} + 2ab + b^{2}\) It this what you are saying? 2) I GIVE you the answer. \(a^{2} + 2ab + b^{2}\) and simply ask you to fill in the third term, \(a^{2} + 2ab\;+\;...\) and you say you don't understand? I am not allowed to believe this. This is extra credit. You have to THINK! What's the point of extra credit if someone just gives it to you?

OpenStudy (candycane):

Just getting it x=-3^2-14

OpenStudy (candycane):

I am not saying the I never multiplied two binomails. I am not asking you to GIVE me the answer, I am asking you to please make it a little more simpler as I am not familar with the equation in the form you have put it let alone how to get to the third term. Believe what you may but I never asked for you to give me the answer straight out

OpenStudy (tkhunny):

Alas. No extra credit for you. If I show you, will you put my name on it and decline the extra credit? I cannot make it simpler. Why do you think it can be made simpler? It is what it is. It is exactly as complicated as it is. It isn't simpler. Okay, I'll bite. Let's try a couple backwards. Please expand (multiple the binomial by itself). \((x+5)^{2}\) \((x-2)^{2}\)